OpenStudy (anonymous):

Automobile Production, The number N of cars produced at a certain factory in 1 day after t hours of operation is given by N(t) = 100t – 5t2, 0 ≤ t ≤ 10. If the cost C (in dollars) of producing N cars is C(N) = 15,000 + 8000N, find the cost C as a function of the time t of operation of the factory.

OpenStudy (anonymous):

what will u get if you replaced the N in the C(N) formula with 100t – 5t2 ?

OpenStudy (anonymous):

@gwenisa r u there?

OpenStudy (anonymous):

C(t) = 15000 + 8000( put here what u have for N(t) ) then expand the answer, can u do that?

OpenStudy (anonymous):

sorry yes i an here

OpenStudy (anonymous):

can you continue from here? you have the answer above , 90% of it, just got to do the remaining 10%, can you?

OpenStudy (anonymous):

it's puzzling

OpenStudy (anonymous):

I just need to know how to set the equation down

OpenStudy (anonymous):

C(t) = 15000 + 8000 (100t - 5t2) C(t) = 15000 + 800000t - 40000t2

OpenStudy (anonymous):

what grade are you in? can you understand the above?

OpenStudy (anonymous):

10th

OpenStudy (anonymous):

then I think the above should be straight forward, right? or do u need me to explain it further?

OpenStudy (anonymous):

i understand the multiplication, but please explain further, thanks

OpenStudy (anonymous):

which part(s) is/are ambiguous for you? can u point it out please?

OpenStudy (anonymous):

to find the cost C

OpenStudy (anonymous):

find the cost C as a function of the time t of operation of the factory ? this part ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok forget about the above and I'll give u a simpler example, just bear with me ok?

OpenStudy (anonymous):

I can make 5 donuts in t hours, ok? so the function for how many donuts I can make in terms of t hours (time that it takes me to make them ) is: N = 5 t after 1 hour I've made 5 donuts after 2 hours I've made 5*2 = 10 donuts after 7 hours I've made 5*7 = 35 donuts is the above clear?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok now to the next part, btw the above equation can be written as is N(t) = 5*t

OpenStudy (anonymous):

The cost C for me to make N number of donuts can be given by this formula C = $2 + 3N in dollars $2 are initial costs for gas burner running all the time so cost of 1 donut = 2+3 = $5 2 donuts ==> 2 + 6 = $8 5 donuts ==> 2 + 15 = $17 do u understand this part?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now they want you to tell them the function of the Cost in terms of the time, you have the formula for the Cost C(N) = 2 + 3N and you have the formula of N in terms of time N(t) = 5t where t is in hours Then all I have to do is find C(t) by replacing each N in the C(N) formula by its N(t) equivalent, i.e. C(t) = 2 + 3 (N) but N = 5t, so C(t) = 2 + 3 (5t) = 2 + 15t this means the cost of making dounts after 1 hour is 2+15 = $17 after 2 hours ==> 2 + 30 = $32 and so on.. how does that sound?

OpenStudy (anonymous):

:) I think i understand what you are saying

OpenStudy (anonymous):

their example is a bit more complex because it got t squared and higher initial costs my example is silly but it conveys the same point cheers

OpenStudy (anonymous):

Thank you very much i think i understand your explanation better. Thanks for your time.

OpenStudy (anonymous):

that's ok, yw